# Plum: Prompt Learning using Metaheuristic

### v2

@@ 2.1 Prompt Learning
Prompt-based learning is an effective approach that harnesses the power of large language models (LLMs) to facilitate downstream tasks by extracting relevant knowledge. This approach offers several advantages, notably in terms of computational efficiency, as only a small set of parameters needs to be optimized, in contrast to traditional methods that require fine-tuning the entire LLM for each task. Prompting methods can be categorized into discrete prompts (Davison et al., 2019; Wallace et al., 2019; Jiang et al., 2020; Shin et al., 2020; Haviv et al., 2021; Yuan et al., 2021; Gao et al., 2021; Ben-David et al., 2022; Su et al., 2022a; Diao et al., 2022; Datta et al., 2023) and continuous prompts (Hambardzumyan et al., 2021; Zhong et al., 2021; Han et al., 2021; Li and Liang, 2021; Qin and Eisner, 2021; Liu et al., 2021a) based on the format of prompts. Discrete prompts are typically represented as token sequences or natural language phrases, while continuous prompts are designed as sequences of vectors. Significant advancements have been made in leveraging large language models (LLMs) for reasoning tasks using chain-of-thought prompting techniques. Notable progress is evidenced by the research of  Wei et al. (2022a); Wang et al. (2022); Zhou et al. (2023); Zhang et al. (2022); Shum et al. (2023); Diao et al. (2023), in which well-structured prompts guide the models to showcase superior reasoning capabilities, surpassing previous benchmarks. Further studies in prompt attacks (Yu et al., 2023; Zhu et al., 2023) have also utilized several of these advancements to improve the robustness of modern LLMs. However, most of the prior studies on prompting have been limited to a white-box setting, requiring access to all parameters of a large language model. The reliance on white-box optimization methods has become less practical in light of the widespread use of closed-source black-box models.
@@ 4.1 Plum for White-Box Prompt Learning
To make the comparison fair for all algorithms, we follow the same editing operators as GrIPS (Prasad et al., 2022), which generally generates discrete prompts by performing phrase-level operations and selecting the best ones. We utilize GPT2-large (Radford et al., 2019) for the backbone model of prompt learning. To accommodate computational resource limitations, we set batch size to $1$ and time-out threshold of searching process to $45$ minutes for all methods. All methods are given a task-agnostic instruction: You will be given a task. Read and understand the task carefully, and appropriately answer [list_of_labels]. Here [list_of_labels] serves as a placeholder to be replaced by an actual list of labels.
@@ Appendix A Discovered Prompt Patterns
Table 5: Plum-HS augmented prompts for stable-diffusion-2-1
Topic
Original Image
Resultant Image
Original Prompt
Augmented Prompt
Cat in a library
A cat prowling in a library at night, books and shadows, silent observer.
A Is that true? Not at all cat prowling in shadows And books shadows,
Moonlight bathes blossom
Moonlight bathes blossom grove, petals shimmer in silvery luminescence
Moonlight bathes blossom silvery There is shimmer in petals
Sunflower
Sunflower field at sunset, golden hour, endless summer.
Sunflower endless field sunset, at sunset,
Moonlight bathes blossom
Moonlight illuminating at night, distant city lights.
Moonlight illuminating at night distant city They have lights
Glowing Mushroom Forest
Illuminated fungi cast ethereal light in dense woodland
Illuminated fungi cast It’s something The and It is ethereal in dense
@@ A Is that true? Not at all cat prowling in shadows And books shadows,
Tasks
Searched Prompts
GSM8K
Initial: $\cdots$ If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? $\cdots$
@@
ASDiv
Initial: $\cdots$ Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total? $\cdots$
@@
AQuA
Initial: $\cdots$ How many keystrokes are needed to type the numbers from 1 to 500? Answer Choices: (a) 1156 (b) 1392 (c) 1480 (d) 1562 (e) 1788$\cdots$
@@
CSQA
Initial: $\cdots$ What do people use to absorb extra ink from a fountain pen? Answer Choices: (a) shirt pocket (b) calligrapher’s hand (c) inkwell (d) desk drawer (e) blotter $\cdots$
@@
StrategyQA
Initial: $\cdots$ Yes or no: Hydrogen’s atomic number squared exceeds number of Spice Girls? $\cdots$
@@ B.1 Plum for White-Box Prompt Learning
More experimental results on Phi-2 [Javaheripi and Bubeck, 2024] and TinyLlama [Zhang et al., 2024] is shown in Table 7, where we still observe on-par or better performance when compare Plum with baseline methods.
@@ C.6 Harmony Search
31:         elsenormal-$\triangleright$ A largely different random segment (pitch)

### v3

@@ 2.1 Prompt Learning
Prompt-based learning is an effective approach that harnesses the power of large language models (LLMs) to facilitate downstream tasks by extracting relevant knowledge. This approach offers several advantages, notably in terms of computational efficiency, as only a small set of parameters needs to be optimized, in contrast to traditional methods that require fine-tuning the entire LLM for each task. Prompting methods can be categorized into discrete prompts (Davison et al., 2019; Wallace et al., 2019; Jiang et al., 2020; Shin et al., 2020; Haviv et al., 2021; Yuan et al., 2021; Gao et al., 2021; Ben-David et al., 2022; Su et al., 2022a; Diao et al., 2022; Datta et al., 2023; Pan et al., 2024) and continuous prompts (Hambardzumyan et al., 2021; Zhong et al., 2021; Han et al., 2021; Li and Liang, 2021; Qin and Eisner, 2021; Liu et al., 2021a) based on the format of prompts. Discrete prompts are typically represented as token sequences or natural language phrases, while continuous prompts are designed as sequences of vectors. Significant advancements have been made in leveraging LLMs for reasoning tasks using chain-of-thought prompting techniques. Notable progress is evidenced by the research of  Wei et al. (2022a); Wang et al. (2022); Zhou et al. (2023); Zhang et al. (2022); Shum et al. (2023); Diao et al. (2023), in which well-structured prompts guide the models to showcase superior reasoning capabilities, surpassing previous benchmarks. Further studies in prompt attacks (Yu et al., 2023; Zhu et al., 2023) have also utilized several of these advancements to improve the robustness of modern LLMs. However, most of the prior studies on prompting have been limited to a white-box setting, requiring access to all parameters of a large language model. The reliance on white-box optimization methods has become less practical in light of the widespread use of closed-source black-box models.
@@ 4.1 Plum for White-Box Prompt Learning
  • APO (Pryzant et al., 2023): automatically improve the initial prompt using beam search guided by textual “gradient”.
@@
To make the comparison fair for all algorithms, we follow the same editing operators as GrIPS (Prasad et al., 2022), which generally generates discrete prompts by performing phrase-level operations and selecting the best ones. We utilize GPT2-large (Radford et al., 2019) for the backbone model of prompt learning. To accommodate computational resource limitations, we set the batch size to $1$ and the time-out threshold of the searching process to $45$ minutes for all methods. All methods are given a task-agnostic instruction: You will be given a task. Read and understand the task carefully, and appropriately answer [list_of_labels]. Here [list_of_labels] serves as a placeholder to be replaced by an actual list of labels.
@@ Results
 We incorporate error bars for all reported results in this paper to illustrate the variability, where $\pm$ represents a plus or minus the standard deviation.
@@ 4.3 Plum for Prompt Discovery
5 Discussion
Plum with LLM-driven Mutators
LLM-based prompt tuning methods are emerging as one of the most competitive approaches in prompt learning. Recent works like Pryzant et al. (2023); Guo et al. (2023) have shown significant performance improvements through utilizing the LLM-driven prompt mutators, highlighting the success of such approaches. Compared with them, Plum offers a complementary and orthogonal paradigm. For example, by incorporating the LLM-based operators in Plum algorithms, our Chain-of-Thought task experiments in Appendix B.2 demonstrate effectiveness in generating high-quality, interpretable few-shot-CoT prompts for reasoning tasks (as shown in Table 6). This suggests that LLMs can be a powerful tool for prompt exploration within the Plum paradigm, particularly when computational resources are abundant.
Strengths of Different Plum Algorithms
In both white-box and black-box settings, our experiments demonstrate that Plum-HS achieves significant performance gains on prompt tuning tasks against baselines and the other Plum algorithms. This is because it takes into account the composability of different prompt segments, which can be seen as an extension of evolutionary algorithms. Notably, in tasks like image generation (detailed in Section 4.3), we observe that updates to different segments can occur in parallel, highlighting this approach’s efficiency.
On the other hand, while two baseline methods may have superior performance over most other Plum algorithms, those metaheuristics possess nice properties that can be beneficial for the community.
  • Plum-HC stands out for its simplicity and efficiency among Plum family. It shares the same form of GrIPS (Prasad et al., 2022) and is one of the most commonly adopted paradigms in LLM-based methods. Plum-HC can serve as a foundational algorithm for developing new mutators. Its straightforward nature allows for easier debugging and comparison when developing more complex algorithms in the Plum family.
  • For Plum-SA, it boasts a quite simple form, which only differs one line from Plum-HC. Crucially, it guarantees to find the global optimum with the infinite time horizon. This characteristic makes it valuable for future theoretical analyses of discrete prompt learning.
  • Plum-GA-M and Plum-GA-C, the genetic algorithm, can be considered one of the most widely adopted algorithms in evolutionary search, whose results can be useful for future comparisons on the same tasks for other researchers. Additionally, our experiments also demonstrate that these algorithms achieve superior performance compared to baseline approaches.
  • For Plum-TS, although its performance is not the best, it achieves on-par performance as GrIPS in  4 times fewer API calls (as shown in Table 3). This advantage makes Plum-TS an extremely favorable choice for applications with limited budgets or API call quotas.
@@ Appendix A Discovered Prompt Patterns
Topic Original Image Resultant Image Original Prompt Augmented Prompt
Cat in a library [Uncaptioned image] [Uncaptioned image] A cat prowling in a library at night, books and shadows, silent observer. A Is that true? Not at all cat prowling in shadows And books shadows,
Moonlight bathes blossom [Uncaptioned image] [Uncaptioned image] Moonlight bathes blossom grove, petals shimmer in silvery luminescence Moonlight bathes blossom silvery There is shimmer in petals
Sunflower [Uncaptioned image] [Uncaptioned image] Sunflower field at sunset, golden hour, endless summer. Sunflower endless field sunset, at sunset,
Moonlight bathes blossom [Uncaptioned image] [Uncaptioned image] Moonlight illuminating at night, distant city lights. Moonlight illuminating at night distant city They have lights
Glowing Mushroom Forest [Uncaptioned image] [Uncaptioned image] Illuminated fungi cast ethereal light in dense woodland Illuminated fungi cast It’s something The and It is ethereal in dense
@@
Tasks Searched Prompts
GSM8K Initial: $\cdots$ If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot? $\cdots$
@@
ASDiv Initial: $\cdots$ Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total? $\cdots$
@@
AQuA Initial: $\cdots$ How many keystrokes are needed to type the numbers from 1 to 500? Answer Choices: (a) 1156 (b) 1392 (c) 1480 (d) 1562 (e) 1788$\cdots$
@@
CSQA Initial: $\cdots$ What do people use to absorb extra ink from a fountain pen? Answer Choices: (a) shirt pocket (b) calligrapher’s hand (c) inkwell (d) desk drawer (e) blotter $\cdots$
@@
StrategyQA Initial: $\cdots$ Yes or no: Hydrogen’s atomic number squared exceeds number of Spice Girls? $\cdots$
@@ B.1 Plum for White-Box Prompt Learning
More experimental results on Phi-2 [Javaheripi and Bubeck, 2024] and TinyLlama [Zhang et al., 2024] are shown in Table 7, where we still observe on-par or better performance when compare Plum with baseline methods. Additionally, Figure 2 illustrates the performance (accuracy) of Plum-HS on the Natural-Instructions dataset subtasks as the number of iterations increases, highlighting the effectiveness and efficiency of our proposed paradigm.
@@
(a) Task 019
(b) Task 021
(c) Task 022
(d) Task 050
(e) Task 069
(f) Task 137
(g) Task 139
(h) Task 195
Figure 2: Performance of Plum-HS across different iterations on subtasks of Natural-Instructions dataset.
@@ C.6 Harmony Search
31:         $\triangleright$ A largely different random segment (pitch)